What is the probability that a coin will turn up heads twice in 6 tosses of the coin?
Every time a coin is tossed there is a 50 / 50 chances of it coming up heads. There is no rule that says tossing it 100 or 6 times will change this.
Each outcome has a probability of 0.05
What is the probability you would get a 2 on both the number cube and the spinner?
The answer depends on how many sides the spinner has.
What is the probability of drawing a two or a three from a deck of cards?
52 cards in a pack.
4 two's, 4 three's, hence 8 chances of drawing one of the desired cards out of the 52.
8/52 = 2/13
What is the probability that two dice add up to five with one toss?
With fair regular dice, the answer is 4/36 = 1/9
Suppose the random variable W represents the weight, and assume that W is Normally distributed with the given mean and standard deviation.
Then Prob(140 < W < 220) = Prob[(140 - 145)/31 < Z < (220 - 145)/31]
where Z has a standard Normal distribution.
that is, Prob(-0.1613 < Z < 2.4194) = 1 - [Prob(-0.1613 < Z) + Prob(Z < 2.4194)
= 1 - (0.4359 + 0.0078) = 0.5563
So percentage of women = 100*0.5563 = 56% approx.
What IS the probability of drawing an ace or a ten from a standard deck of cards?
In a deck of 52 cards, there are 4 chances to get an ace and 4 chances to get a ten so there are 8 chances to get an ace or a ten
So the probability is 8/52=2/13=0.15
What is the probability of obtaining exactly three heads in seven throws with a single coin?
(7!/(4!*3!))*(1/2)^3 * (1/2)^4 = 35*(1/2)^7 which is about 0.2734375
Every time you flip a coin it has a 50% probability that it will land on either heads or tails. You would expect to get heads about half the time and tails about half the time. What actually happens could be different from what is expected. You could get heads every time, or tails every time. Or you could get tails 75% of the time and heads 25% of the time. however, your results appear to be what you would expect, approximately 50% heads and 50% tails. You got 16 heads and 14 tails. Your percentage of heads is 16/30 x 100= 53.33... %. Your percentage of tails is 14/30 x 100 = 46.66... %.
What is the probability of obtaining a number less than three at the throw of a dice?
There are 6 values, 1 - 6, on the die. Less than 3 is a 1 or 2 roll. So, probability of less than 3 is 2/6 or 1/3 or 33.33%.
How many 5 card poker hands contain only Diamonds?
There are 1287 possible 5 card poker hands that contain only diamonds, i.e. a diamond flush.
This is 13! / (13-5)! / 5!
Since there are 2598960 possible 5 card poker hands, 52! / (52-5)! / 5!, the probability of drawing a diamond flush is 1287 in 2598960, or about 0.0004952.
What is the probability of rolling a sum of 5 when rolling two 6 sided dice?
The answer is 1/9 because the possibilities are (1,4), (2,3), (3,2), and (4,1). 4 outcomes out of a total 36 possible: 4/36 = 1/9 = 11.11% chance
Yes, it is an independent event. However, if you did not replace the card it would not be an independent event, since you don't have an equal chance of pulling any particular card.
The outcome space is TT, TH, HT, HH: 4 outcomes in all.
{One head} = {TH, HT} with prob = 2/4 = 1/2
{No heads} = {TT} with prob 1/4
So E = 2*Pr(one H) - 4*Pr(no H) = 2*(1/2)-4*(1/4) = 1 - 1 = 0
What are the chances of your house getting burgled?
It would depend on the location and crime rate.
Check the local crime statistics for the area.
In the right order, it's (1/4) x (1/3) x (1/2) = 1/24
In either order it's twice as much: 1/12.
The probability of an event must be a number in the interval [0, 1].
IS a Joker included as a face card?
it could go either way because in games you don't play with it, but in trash some people say it is a free card.
1/48 if there aren't jokers
What is the difference between a biased sample and a random sample?
a biased sample is chosen because it is liked more (looks prettier, more popular etc.) and a random one has no prefeerence hope i helped!!!! :p
What does the denominator represent in probability?
It represents the total number of outcomes of a trial divided by the highest common factor of the number of favourable outcomes and the total number of outcomes.